OpenStudy (anonymous):

plz help use a transformation to solve the equation. d – (–17) = –14 A. d = –31 B. d = –3 C. d = 3 D. d = 31

OpenStudy (anonymous):

d = –31

OpenStudy (anonymous):

-31

OpenStudy (anonymous):

@travolta1988 please don't give direct answers

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

d = -14 + (-17) = -14-17 = -(14 +17) = -31

OpenStudy (anonymous):

 solve it @surviver101